# Application of the Lovász-Schrijver Lift-and-Project Operator to Compact Stable Set Integer Programs

### v2

@@ 4.2 Relaxation $N_{+}(\operatorname{NOD}(G,r))$
$x_{j}\cdot\left(\sum_{h\in\Gamma(i)}x_{h}+r_{i}x_{i}\leq r_{i}\right)\rightarrow$ $\sum_{h\in\Gamma(i)}X_{ih}\leq 0,$ if $j=i$ (24)
$x_{j}\cdot\left(\sum_{h\in\Gamma(i)}x_{h}+r_{i}x_{i}\leq r_{i}\right)\rightarrow$ $(1-r_{i})x_{j}+\sum_{h\in\Gamma(i)\setminus j}X_{jh}+r_{i}X_{ij}\leq 0,$ if $j\in\Gamma(i)$ (25)
$x_{j}\cdot\left(\sum_{h\in\Gamma(i)}x_{h}+r_{i}x_{i}\leq r_{i}\right)\rightarrow$ $\sum_{h\in\Gamma(i)}X_{jh}+r_{i}X_{ij}-r_{i}x_{j}\leq 0,$ if $j\not\in\Gamma(i)$ (26)
$(1-x_{j})\cdot\left(\sum_{h\in\Gamma(i)}x_{h}+r_{i}x_{i}\leq r_{i}\right)\rightarrow$ $\sum_{h\in\Gamma(i)}(x_{h}-X_{ih})+r_{i}x_{i}\leq r_{i},$ if $j=i$ (27)
$(1-x_{j})\cdot\left(\sum_{h\in\Gamma(i)}x_{h}+r_{i}x_{i}\leq r_{i}\right)\rightarrow$ $\sum_{h\in\Gamma(i)\setminus j}(x_{h}-X_{jh})+r_{i}x_{i}+r_{i}x_{j}-r_{i}X_{ij}\leq r_{i},$ if $j\in\Gamma(i)$ (28)
$(1-x_{j})\cdot\left(\sum_{h\in\Gamma(i)}x_{h}+r_{i}x_{i}\leq r_{i}\right)\rightarrow$ $\sum_{h\in\Gamma(i)}(x_{h}-X_{jh})+r_{i}x_{i}+r_{i}x_{j}-r_{i}X_{ij}\leq r_{i},$ if $j\not\in\Gamma(i)$ (29)
@@
    Constraints (24) imply $X_{ij}=0$ for all $\{i,j\}\in E$; this is a relevant fact if one recalls that edge inequalities are not implied by nodal inequalities (9);
@@
    $X_{ij}=0$, for all $\{i,j\}\in E$, $X_{ij}\geq 0$ for all $\{i,j\}\in\bar{E}$ and constraints (26) imply inequalities (15) and (16);
@@
    $X_{ij}=0$, for all $\{i,j\}\in E$, inequalities (15) and (16) and constraints (28) imply edge inequalities (and (17) for $\{i,j\}\in E$);
@@
    By $X_{ij}=0$, for all $\{i,j\}\in E$, (27) boils down to the original nodal inequality (9) and can be dropped, since Remark 1 can be applied to both pairs (25), (28) and/or (26), (29);
@@
    $X_{ij}\geq 0$, for all $\{i,j\}\in\bar{E}$, inequalities (15) and (16) and constraints (29) imply (17) for $\{i,j\}\in\bar{E}$.
@@
$\begin{split}{M_{+}(\operatorname{NOD}(G,r))}=\left\{\begin{pmatrix}1&x^{\top}\\ x&X\end{pmatrix}\succeq 0:\right.&\\ (1-r_{i})x_{j}+\sum_{h\in\Gamma(i)\setminus j}X_{jh}\leq 0,&\quad i\in V,j\in\Gamma(i),\\ \sum_{h\in\Gamma(i)}X_{jh}+r_{i}X_{ij}-r_{i}x_{j}\leq 0,&\quad i\in V,j\not\in\Gamma(i),\\ \sum_{h\in\Gamma(i)\setminus j}(x_{h}-X_{hj})+r_{i}x_{i}+r_{i}x_{j}\leq r_{i},&\quad i\in V,j\in\Gamma(i),\\ \sum_{h\in\Gamma(i)}(x_{h}-X_{hj})+r_{i}x_{i}+r_{i}x_{j}-r_{i}X_{ij}\leq r_{i},&\quad i\in V,j\not\in\Gamma(i),\\ x_{i}=X_{ii},&\quad i\in V,\\ X_{ij}=0,&\quad\{i,j\}\in E,\\ X_{ij}\geq 0,&\quad\{i,j\}\in{\bar{E}},\\ x_{i}\geq 0,&\quad i\in V\bigg{\}}.\end{split}$ (30)
Then, $M_{+}(\operatorname{NOD}(G,r))$ can be obtained by adding linear inequalities (25), (26), (28), (29) and $X_{ij}\geq 0$, for $\{i,j\}\in\bar{E}$, to (SDP-$\theta$) (it is straightforward to observe that inequalities (25) are implied by Mc-Cormick inequalities for $r_{i}=\Gamma$). Therefore, its projection $N_{+}(\operatorname{NOD}(G,r))$ onto the $x$-space is at least as strong as the Schrijver relaxation:
@@
Which also implies:
@@ 4.3 Comparison among relaxations
Consider the odd-hole with seven vertices of Figure (2(b)). According to [43], $N_{+}(\operatorname{FRAC}(G))$ satisfies the facet-defining inequality $\sum_{i=1}^{8}x_{i}\leq 3$ which implies $\lambda(G)=3$. Differently, the optimal value from $N_{+}(\operatorname{NOD}(G,r))$ is $\nu(G,\theta)=\nu(G,\alpha)=3.317$ corresponding to solution $(0.47395,\ldots,0.47395)$, which therefore falls outside $N_{+}(\operatorname{FRAC}(G))$. However, one can find a feasible solution to $N_{+}(\operatorname{FRAC}(G))$ which is outside $N_{+}(\operatorname{NOD}(G,r))$ as well. Consider the $(10,3)$-antiweb graph $G$ of Figure (2(a)). Here, $\nu(G,\theta)=\nu(G,\alpha)=\alpha(G)=3$. On the contrary, optimizing over $N_{+}(\operatorname{FRAC}(G))$ we obtain solution
@@
This example also shows that the odd-hole inequalities are not implied by $N_{+}(\operatorname{NOD}(G,r))$ even with $r=\alpha$. This appears as a weakness indicator of $N_{+}(\operatorname{NOD}(G,r))$ if one recalls that odd-hole inequalities are contained in the first Chvàtal closure of the edge formulation 1 [53]. The overall theoretical picture is summarized in Figure 3, and the above results seem to suggest that $N_{+}(\operatorname{QSTAB}(G))$ (as well as $N_{+}(\operatorname{FRAC}(G))$) should be stronger than $N_{+}(\operatorname{NOD}(G,r))$. Interestingly, the computational analysis will reveal significant supremacy of the latter as graphs get dense, for $r\in\{\theta,\alpha\}$.
Symbol Relaxation
@@ 5 Implementation
[…]
@@
[…]
@@
The function m_plus_lifting() in SDPLifting.py carries out Steps 1 and 2 of the procedure outlined in Section 2.1. Initially, it parses the polytope $P$, provided as input through a standard .lp format text file. Then, it construct the symmetric matrices corresponding to constraints (4), (5) and (7). Finally, the model is exported in .mat file format, ready to be loaded into MATLAB and solved using SDPNAL+.
@@
and $m\in\mathbb{N}$, $C,A_{1},\ldots,A_{m}\in{\cal S}_{n+1}$, $b\in\mathbb{R}^{m}$. In a preliminary experiment we observed that SDPNAL+ converges quickly on (SDP-$\theta^{+}$) while it requires very large (and practically unacceptable) convergence times on $M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))$ and $M_{+}(\operatorname{NOD}(G,r))$ relaxations. This is due to the number of linear constraints in the former, and their density in the latter. However, we proved that $M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))$ and $M_{+}(\operatorname{NOD}(G,r))$ are contained in (SDP-$\theta^{+}$), thus one can resort to a classical cutting plane method to optimize over them. The cutting plane algorithm starts optimizing over (SDP-$\theta^{+}$). At the $k$-th iteration, given the optimal solution $Y^{k}\in{\cal S}^{+}_{n+1}$, one can compute
@@
[…]
@@ 6 Computational Experiments
    Can relaxations based on operator $M_{+}(\cdot)$ provide upper bounds for $\alpha(G)$ that are significantly stronger than $\theta(G)$?
@@
As a byproduct, we document the quality of the bound provided by $M_{+}(\operatorname{NOD}(G,\theta))$, that is, the strongest bound which can be computed in polynomial time in our hierarchy.
Experiments are run on a computer with Intel(R) Xeon(R) CPU E5-2698 v4 @ 2.20GHz, 256GB RAM, under OS Ubuntu 16.04.7 LTS (GNU/Linux 4.15.0-128-generic x86_64). After preliminary tuning experiments, the parameters of Algorithm 1 have been set to $\epsilon=10^{-3},\delta=10^{-1}$ and $c=1000$ with 7200 seconds time limit (excluding formulation building time). The precision of the SDP solver SDPNAL+ is set to $10^{-6}$. The precision of ADMMsolver.py for evaluating coefficients $\theta_{i}$ is set to $10^{-4}$. Denoting by $\operatorname{opt}(\operatorname{R})$ the optimal (with the given precision) value of relaxation $\operatorname{R}$ computed by Algorithm 1 at termination, tables below report, for each relaxation, four statistics: the percentage gap to $\alpha(G)$, that is, $\frac{\operatorname{opt}(\operatorname{R})-\alpha(G)}{\alpha(G)}\%$; the number of cutting plane iterations and the corresponding CPU time; the number of added cuts. The latter is detailed for each class of inequalities and a class is not reported if none of its members has ever been detected as violated.

### v3

@@ 4.2 Relaxation $N_{+}(\operatorname{NOD}(G,r))$
$x_{j}\cdot\left(\sum_{k\in\Gamma(i)}x_{k}+r_{i}x_{i}\leq r_{i}\right)\rightarrow$
$\rightarrow$ $\sum_{k\in\Gamma(i)}X_{ik}\leq 0,$ if $j=i$ (15a)
$\rightarrow$ $(1-r_{i})x_{j}+\sum_{k\in\Gamma(i)\setminus j}X_{jk}+r_{i}X_{ij}\leq 0,$ if $j\in\Gamma(i)$ (15b)
$\rightarrow$ $\sum_{k\in\Gamma(i)}X_{jk}+r_{i}X_{ij}-r_{i}x_{j}\leq 0,$ if $j\not\in\Gamma(i)$ (15c)
$(1-x_{j})\cdot\left(\sum_{k\in\Gamma(i)}x_{k}+r_{i}x_{i}\leq r_{i}\right)\rightarrow$
$\rightarrow$ $\sum_{k\in\Gamma(i)}(x_{k}-X_{ik})+r_{i}x_{i}\leq r_{i},$ if $j=i$ (16a)
$\rightarrow$ $\sum_{k\in\Gamma(i)\setminus j}(x_{k}-X_{jk})+r_{i}x_{i}+r_{i}x_{j}-r_{i}X_{ij}\leq r_{i},$ if $j\in\Gamma(i)$ (16b)
$\rightarrow$ $\sum_{k\in\Gamma(i)}(x_{k}-X_{jk})+r_{i}x_{i}+r_{i}x_{j}-r_{i}X_{ij}\leq r_{i},$ if $j\not\in\Gamma(i)$ (16c)
@@
    Constraints (15a) imply $X_{ij}=0$ for all $\{i,j\}\in E$; this is a relevant fact if one recalls that edge inequalities are not implied by nodal inequalities (9);
@@
    $X_{ij}=0$, for all $\{i,j\}\in E$, $X_{ij}\geq 0$ for all $\{i,j\}\in\bar{E}$ and constraints (15c) imply inequalities (12b) and (12c);
@@
    $X_{ij}=0$, for all $\{i,j\}\in E$, inequalities (12b) and (12c) and constraints (16b) imply edge inequalities (and (12d) for $\{i,j\}\in E$);
@@
    By $X_{ij}=0$, for all $\{i,j\}\in E$, (16a) boils down to the original nodal inequality (9) and can be dropped, since Remark 1 can be applied to both pairs (15b), (16b) and (15c), (16c);
@@
    $X_{ij}\geq 0$, for all $\{i,j\}\in\bar{E}$, inequalities (12b) and (12c) and constraints (16c) imply (12d) for $\{i,j\}\in\bar{E}$.
@@
$\begin{split}{M_{+}(\operatorname{NOD}(G,r))}=\left\{\begin{pmatrix}1&x^{\top}\\ x&X\end{pmatrix}\succeq 0:\right.&\\ (1-r_{i})x_{j}+\sum_{k\in\Gamma(i)\setminus j}X_{jk}\leq 0,&\quad i\in V,j\in\Gamma(i),\\ \sum_{k\in\Gamma(i)}X_{jk}+r_{i}X_{ij}-r_{i}x_{j}\leq 0,&\quad i\in V,j\not\in\Gamma(i),\\ \sum_{k\in\Gamma(i)\setminus j}(x_{k}-X_{jk})+r_{i}x_{i}+r_{i}x_{j}\leq r_{i},&\quad i\in V,j\in\Gamma(i),\\ \sum_{k\in\Gamma(i)}(x_{k}-X_{jk})+r_{i}x_{i}+r_{i}x_{j}-r_{i}X_{ij}\leq r_{i},&\quad i\in V,j\not\in\Gamma(i),\\ x_{i}=X_{ii},&\quad i\in V,\\ X_{ij}=0,&\quad\{i,j\}\in E,\\ X_{ij}\geq 0,&\quad\{i,j\}\in{\bar{E}},\\ x_{i}\geq 0,&\quad i\in V\bigg\}.\end{split}$
Then, $M_{+}(\operatorname{NOD}(G,r))$ can be obtained by adding linear inequalities (15b), (15c), (16b), (16c) and $X_{ij}\geq 0$, for $\{i,j\}\in\bar{E}$, to (3.2.2) (it is straightforward to observe that inequalities (15b) are implied by Mc-Cormick inequalities for $r=\Gamma$). Therefore, its projection $N_{+}(\operatorname{NOD}(G,r))$ onto the $x$-space is at least as strong as the Schrijver relaxation.
@@
This result also implies
@@ 4.3 Comparison among relaxations
Consider the odd-hole with seven vertices of Figure (2(b)). According to [43], $N_{+}(\operatorname{FRAC}(G))$ satisfies the facet-defining inequality $\sum_{i\in V}x_{i}\leq 3$ which implies $\lambda(G)=3$. Differently, the optimal value from $N_{+}(\operatorname{NOD}(G,r))$ is $\nu(G,\theta)=\nu(G,\alpha)=3.317$ corresponding to solution $(0.47395,\ldots,0.47395)$, which therefore falls outside $N_{+}(\operatorname{FRAC}(G))$. However, one can find a feasible solution to $N_{+}(\operatorname{FRAC}(G))$ which is outside $N_{+}(\operatorname{NOD}(G,r))$ as well. Consider the $(10,3)$-antiweb graph $G$ of Figure (2(a)). Here, $\nu(G,\theta)=\nu(G,\alpha)=\alpha(G)=3$. On the contrary, optimizing over $N_{+}(\operatorname{FRAC}(G))$ we obtain the solution
@@
This example also shows that the odd-hole inequalities are not implied by $N_{+}(\operatorname{NOD}(G,r))$ even with $r=\alpha$. This appears as a weakness indicator of $N_{+}(\operatorname{NOD}(G,r))$ if one recalls that odd-hole inequalities are contained in the first Chvàtal closure of the formulation based on edge inequalities (1) [53]. The overall theoretical picture is summarized in Figure 3 . The results related to the implication of odd-hole and antiweb inequalities seem to suggest that $N_{+}(\operatorname{QSTAB}(G))$ (as well as $N_{+}(\operatorname{FRAC}(G))$) should be stronger than $N_{+}(\operatorname{NOD}(G,r))$. Interestingly, the computational analysis will reveal significant supremacy of the latter as graphs get dense, for $r\in\{\theta,\alpha\}$.
Optimal value Relaxation
@@ 5 Implementation
Under mild assumptions, semidefinite programs can be solved in polynomial time up to any arbitrary fixed precision [51]. Intuitively, the computational burden mainly depends on the order $n$ of the matrix variable and the number $m$ of linear constraints. However, the density of the latter should also be taken into account as most SDP solvers rely on sparse matrices implementations. Interior-point methods (IPMs) [47, 60, 64] provide a good-precision solution in reasonable time for “small” and “medium” SDPs (both in term of $n$ and $m$). However, they require to store and factorize large Hessian matrices at each iteration, which may become prohibitive for large SDPs. In fact, we preliminarily tested the IPM implemented in the commercial state-of-the-art solver MOSEK [47] and observed that problems with $m\approx 10^{5}$ and $n\approx 200$ are intractable. Although this experiment already provided interesting insights on the strength of the relaxations, testing larger graphs is crucial for a reliable experimental picture. Therefore, we resorted to Alternating Direction Methods of Multipliers (ADMMs), which, as a variant of Augmented Lagrangian methods, is a popular first-order alternative to IPMs and can scale on much larger SDPs at the price of possibly slowing down the convergence to high-precision solutions. For a detailed overview, we refer the reader to [4, 5, 17, 54, 61, 65, 66]. Furthermore, ADMMs allow the efficient handling of bound constraints (e.g., nonnegativity) on the matrix variable via iterative projections on the cone of positive semidefinite matrices. In our case, it enables the computation of $\theta^{+}(G)$ with no substantial additional effort with respect to $\theta(G)$. We selected SDPNAL+ proposed in [61], which represents a state-of-the-art ADMM general-purpose SDP solver implemented in MATLAB equipped with efficient subroutines in C included via .mex files, and has proven to be numerically stable in our experiments.
@@
The $M_{+}(\cdot)$ operator has been applied to four different compact LP relaxations of the SSP: $\operatorname{QSTAB}(G,\mathcal{C})$ and $\operatorname{NOD}(G,r)$ for $r\in\{\Gamma,\theta,\alpha\}$. The construction of these LP models entails different complexity levels. $\operatorname{QSTAB}(G,\mathcal{C})$ requires the computation of an edge clique cover $\mathcal{C}\in\mathcal{Q}$ . This task has been accomplished by the greedy algorithm presented in Letchford et al. [40] and is reported in Algorithm 1. Starting from the empty collection $\mathcal{C}=\emptyset$, the heuristic iteratively adds a maximal clique of $G$ to $\mathcal{C}$, until every edge is covered. At each iteration, the selected maximal clique $C_{k}$ covers at least one new edge not yet covered by cliques in $\mathcal{C}$ (line 8). Consequently, even if it can be shown that the resulting edge clique cover $\mathcal{C}$ may not be minimal, its size is at most $|E|$ and, in practice, comprises significantly fewer cliques [40].
Algorithm 1 Greedy edge clique cover
1:Input: Graph $G=(V,E)$
2:Output: An edge clique cover $\mathcal{C}$ of $G$
3:Initialize $E^{\prime}\leftarrow E,$ $G^{\prime}\leftarrow(V,E^{\prime})$ $\mathcal{C}\leftarrow\emptyset,\ k\leftarrow 0$
4:while $E^{\prime}\not$$\emptyset$ do
5:  $k\leftarrow k+1$
6:  $u\leftarrow\arg\max_{i\in V}\left\{|\Gamma_{G^{\prime}}(i)|\right\}$
7:  $v\leftarrow\arg\max_{i\in\Gamma_{G^{\prime}}(u)}\left\{|\Gamma_{G^{\prime}}(i)|\right\}$ $\triangleright$ $E^{\prime}\not=\emptyset\Rightarrow\Gamma_{G^{\prime}}(u)$ is non-empty
8:  Find a maximal clique $C_{k}$ of $G$ such that $u,v\in C_{k}$
9:  $\mathcal{C}\leftarrow\mathcal{C}\cup\left\{C_{k}\right\}$
10:  $E^{\prime}\leftarrow E^{\prime}\setminus E(C_{k})$
11:end while
Building $\operatorname{NOD}(G,\Gamma)$ is straightforward, as $r_{i}=|\Gamma(i)|$. On the contrary, $\operatorname{NOD}(G,\theta)$ requires the evaluation of $r_{i}=\theta(G[\Gamma(i)])$, for all $i\in V$. This corresponds to solve $|V|$ SDPs, namely, those corresponding to (SDP-$\theta$) on the subgraph induced by $\Gamma(i)$. In practice, the SDPs are completely decomposed and can be solved in parallel. Moreover, $r_{i}$ can be tightened to $\left\lfloor\theta(G[\Gamma(i)])\right\rfloor$ and one can stop the computation as long as the SDP solver returns a valid upper (dual) bound on $\theta(G[\Gamma(i)])$. Several approaches to compute a valid dual bound throughout solver’s iterations and a posteriori have been proposed for ADMM algorithms (see, e.g., [17, 36, 66]). Their specific application in SDP relaxations of the SSP have been implemented in [4, 5]. Then, for each iteration of the ADMM implemented in ADMMsolver.py, a valid dual bound is computed and this, in turn, allows stopping the computation of $\theta$ as soon as a “moderate” precision is achieved. In Table 2, we report the average/max/min CPU times to calculate coefficients $\theta(G[\Gamma(i)])$ on a subset of the instances used in our computational experiments, whose detailed introduction is deferred to Section 6.1. Finally, building $\operatorname{NOD}(G,\alpha)$ is the most challenging case as it requires calculating $\alpha(G[\Gamma(i)])$, for all $i\in V$. Despite the theoretical hardness of determining $\alpha(G[\Gamma(i)])$Letchford et al. [40] showed that in practice, the computational burden to build $\operatorname{NOD}(G,\alpha)$ is, in fact, accessible for large classes of graphs. We refer the interested reader to their discussion; here we mention that even in this case the evaluation of coefficients $r_{i}=\alpha(G[\Gamma(i)])$ can be parallelized and, as in the previous case, we report in Table 2 the average/max/min CPU times to evaluate $\alpha(G[\Gamma(i)])$ via the max-clique solver cliquer [52].
@@
The function m_plus_lifting() in SDPLifting.py carries out Steps 1 and 2 of the procedure outlined in Section 2.1. Initially, it parses the polytope $P$, provided as input through a standard .lp format text file. Then, it constructs the symmetric matrices corresponding to constraints (4), (5) and (7). Finally, the model is exported in .mat file format, ready to be loaded into MATLAB and solved using SDPNAL+.
@@
and $m\in\mathbb{N}$, $C,A_{1},\ldots,A_{m}\in{\cal S}_{n+1}$, $b\in\mathbb{R}^{m}$. In a preliminary experiment we observed that SDPNAL+ converges quickly on (3.2.2) while it requires very large (and practically unacceptable) convergence times on $M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))$ and $M_{+}(\operatorname{NOD}(G,r))$ relaxations. This is due to the number of linear constraints in the former, and their density in the latter. However, we proved that $M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))$ and $M_{+}(\operatorname{NOD}(G,r))$ are contained in (3.2.2), thus one can resort to a classical cutting plane method to optimize over them. The cutting plane algorithm starts optimizing over (3.2.2). At the $k$-th iteration, given the optimal solution $Y^{k}\in{\cal S}^{+}_{n+1}$, one can compute
@@
If $V_{i}(Y^{k})>\varepsilon$, for a “small” $\varepsilon>0$, then the $i$-th constraint is violated by the current solution. Let $c\leq m$ be a positive integer parameter, and let $I_{c}(Y^{k})=\left\{i_{1},i_{2},\ldots,i_{c}\right\}$ denote the set of indexes corresponding to the $c$ most violated constraints. Then, the constraints $\left\{Y\in{\cal S}_{n+1}\ :\ \langle A_{i},Y\rangle\leq b_{i},\ i\in I_{c}(Y^{k})\right\}$ are added to the current SDP relaxation and a new optimal solution $Y^{k+1}$ is computed. The algorithm terminates either when no violated constraints are found or when a tailing off condition is met. The detailed algorithm is reported in Algorithm 2.
@@ 6 Computational Experiments
    Can relaxations based on operator $M_{+}(\cdot)$ provide upper bounds on $\alpha(G)$ that are significantly stronger than $\theta(G)$?
@@
As a byproduct, we document the quality of the bound provided by $M_{+}(\operatorname{NOD}(G,\theta))$, that is, the strongest bound which can be computed in polynomial time in our hierarchy. Finally, we investigate the extent to which the bound $\mu(G,\mathcal{C})$ is sensitive to the choice of $\mathcal{C}$ and, in particular, the degree to which it can be improved with a more refined selection of the edge clique cover.
Experiments are run on a computer with Intel(R) Xeon(R) CPU E5-2698 v4 @ 2.20GHz, 256GB RAM, under OS Ubuntu 16.04.7 LTS (GNU/Linux 4.15.0-128-generic x86_64). After preliminary tuning experiments, the parameters of Algorithm 2 have been set to $\varepsilon=10^{-3},\delta=10^{-1}$ and $c=1000$ with 7200 seconds time limit (excluding formulation building time). The precision of the SDP solver SDPNAL+ is set to $10^{-6}$. The precision of ADMMsolver.py for evaluating coefficients $\theta_{i}$ is set to $10^{-4}$. Denoting by $\operatorname{opt}(\operatorname{R})$ the optimal (with the given precision) value of relaxation $\operatorname{R}$ computed by Algorithm 2 at termination, tables below report, for each relaxation, four statistics: the percentage gap to $\alpha(G)$, that is, $\frac{\operatorname{opt}(\operatorname{R})-\alpha(G)}{\alpha(G)}\%$; the number of cutting plane iterations and the corresponding CPU time; the number of added cuts. The latter is detailed for each class of inequalities and a class is not reported if none of its members has ever been detected as violated.